OpenStudy (anonymous):

Anyone willing to help with this problem? I was just in the hospital and my class ends tonight. I'm so stressed out. A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is 65% salt and Solution B is 90% salt. She wants to obtain 100 ounces of a mixture that is 75% salt. How many ounces of each solution should she use?

OpenStudy (anonymous):

If somebody could help explain how to set the problem up and work with me step by step

OpenStudy (tkhunny):

Equate a single substance. Definition: x = Amount of Solution A Definition: 100 - x = Amount of Solution B You need to see where that came from. Equate Salt x * 0.65 + (100-x) * 0.90 = 100 * 0.75 Equate whatever is not salt. x * (1-0.65) + (100-x) * (1-0.90) = 100 * (1-0.75) x * 0.35 + (100-x) * 0.10 = 100 * 0.25 They don't look like they'll give the same solution, do they? Do it!

OpenStudy (anonymous):

x * 0.65 + (100-x) * 0.90 = 100 * 0.75 one is 60, Let me work out the rest

OpenStudy (anonymous):

I'm confused about the rest of the equations

OpenStudy (anonymous):

The second one is easy. Remember that Amount of B = 100-Amount of A. So once you find A, you only need to plug into that equation to find Amount of B.

OpenStudy (anonymous):

Do I plug in 60?

OpenStudy (anonymous):

Yes. If (Amount of B) = 100 - (Amount of A) and (Amount of A) = 60, simply plug it in to get Amount of B.

OpenStudy (anonymous):

Ok, thank you!

OpenStudy (tkhunny):

"one is 6" -- Where did you get that? Did you actually understand, use, and solve one of the equations?